A bob sled race (down a straight course) is able to enter the starting gate with a speed of 4.0 m/sec and the average acceleration down the hill was 6.2 m/sec^2. If the sled covers the course in 1.35 seconds, how fast was it moving at the bottom of the course?
1
Expert's answer
2010-10-03T23:44:15-0400
It is moving with constant acceleration so we have

v-v_0 = a*t => v=a*t+v_0

so

v=6.2*1.35 + 4 = 12.37 m/s